# Strength Lies in Differences! Towards Effective Non-collaborative Dialogues via Tailored Strategy Planning

### v2

@@ A.2 Evaluation Tasks
Following Bianchi et al. (2024); Deng et al. (2023c), we consider two classic tasks as our evaluation scenarios, including price negotiation He et al. (2018) and charity persuasion Wang et al. (2019). The price negotiation task involves open-ended price negotiations where a buyer influences the seller towards a reasonable price, while the seller aims to maximize their own profit. The charity persuasion task involves asymmetric interactions guided by a persuader who endeavors to persuade the other party to make a charitable donation. Our evaluation is based on these two tasks, requiring the evaluated dialogue agents to take on the roles of buyer and persuader, respectively, in order to achieve their goals. To support our evaluations, we adopt the test dataset of CraigslistBargain He et al. (2018) and PersuasionForGood Wang et al. (2019), making use of their pre-annotated background information to streamline our assessment process. For the negotiation task, the background information includes item details and the desired price of each party. For the persuasion task, it involves determining if the individual being persuaded initially intends to make a donation. These background information serve as specific scenarios for our evaluation.
@@ A.3 Reliability Analysis
Prior to conducting the interactive evaluation, we validate the reliability of using LLMs as user simulators that demonstrate non-collaborative behaviors. Following the approach described in Deng et al. (2023c), we engage 5 human experts in conversations with two groups, including our diverse user simulators and 10 real users across two evaluation tasks. We collect 50 dialogues from each group and evaluate the user responses in both single-turn and multi-turn open-ended conversations. The evaluation focuses on the naturalness and utility of the generated responses in these conversation settings. Naturalness refers to the fluency and human-like nature of the responses, while utility indicates their consistency with the role instructions and non-collaborative behaviors. We employ two annotators to conduct pairwise evaluations by rating "Win/Tie/Lose" between the two samples. As shown in Table 8, the user simulators exhibit a notably superior performance compared to real users, particularly when it comes to the naturalness of responses in multi-turn conversations, which showcases the impressive language generation capabilities inherent in LLMs. Furthermore, even compared with human-annotated dialogues, the GPT3.5-based simulator shows competitive performance. These results validate the reliability of adopting GPT3.5 as the user simulator.
@@ A.4 Interactive Evaluation Protocol
During the evaluation, each dialogue agent must engage with these simulators Deng et al. (2023c). During interactions, the dialogue agent and user simulator alternate in employing strategies in their responses with the ultimate aim of maximizing their own self-interest. The interactions continues until the conversational goal is achieved or the maximum number of turns T (i.e., T is set to 10 for both tasks) is reached. To determine goal achievement, we utilize AI feedback to assess whether the task goal has been reached. Specifically, in price negotiation task, we employ a separate GPT3.5 (i.e., $LLM_{rwd}$) to assess whether both parties have reached a deal. We prompt $LLM_{rwd}$ to generate feedback for the binary question “Have they reached a deal?”. If the output of $LLM_{rwd}$ indicates that both parties have reached an agreement, we consider this as goal achievement. In charity persuasion task, we additionally prompt the user simulator to express his willingness to make a donation at the end of each turn. In particular, we query the user simulator "Would you be interested in donating to Save the Children?". If the feedback is positive, we regard this as goal achievement. Conversely, if the goal is not achieved, the interaction continues.
Due to the subjectivity of the planning outcome as well as the variance of the LLM-generated output, we follow a common practice Wang et al. (2022); Deng et al. (2023c) to alleviate these issues by sampling the decoded sequences l (i.e., l is set to 10 for both tasks) times.
@@ B.2 Baselines Implementation Details
ProCoT: we follow Deng et al. (2023b) and prompt LLM to analyze the dialogue status and plan next strategy, and then generate a response based on the planned strategy. We provide its prompt design in Table 17.
@@
PPDPP: we follow Deng et al. (2023c) and adopt the BERT [^7] https://huggingface.co/google-bert/bert-base-uncased model Devlin et al. (2018) as our external planner. We implement PPDPP based on the training details provided in the original paper. We have made adjustments to the task instructions and role-playing descriptions, adapting them for use in the context of charity persuasion.

### v3

@@ A.1 Building User Simulators
Due to the significant human labor required for real-user evaluations Huang et al. (2023), our experiments utilize user simulators instead.
@@ A.2 Evaluation Tasks
Following Bianchi et al. (2024); Deng et al. (2023e), we consider two classic tasks as our evaluation scenarios, including price negotiation He et al. (2018) and charity persuasion Wang et al. (2019). The price negotiation task involves open-ended price negotiations where a buyer influences the seller towards a reasonable price, while the seller aims to maximize their own profit. The charity persuasion task involves asymmetric interactions guided by a persuader who endeavors to persuade the other party to make a charitable donation. Our evaluation is based on these two tasks, requiring the evaluated dialogue agents to take on the roles of buyer and persuader, respectively, in order to achieve their goals. To support our evaluations, we adopt the test dataset of CraigslistBargain He et al. (2018) and PersuasionForGood Wang et al. (2019), making use of their pre-annotated background information to streamline our assessment process. For the negotiation task, the background information includes item details and the desired price of each party. For the persuasion task, it involves determining if the individual being persuaded initially intends to make a donation. These background information serve as specific scenarios for our evaluation.
@@ A.3 Reliability Analysis
Prior to conducting the interactive evaluation, we validate the reliability of using LLMs as user simulators that demonstrate non-collaborative behaviors. Following the approach described in Deng et al. (2023e), we engage 5 human experts in conversations with two groups, including our diverse user simulators and 10 real users across two evaluation tasks. We collect 50 dialogues from each group and evaluate the user responses in both single-turn and multi-turn open-ended conversations. The evaluation focuses on the naturalness and utility of the generated responses in these conversation settings. Naturalness refers to the fluency and human-like nature of the responses, while utility indicates their consistency with the role instructions and non-collaborative behaviors. We employ two annotators to conduct pairwise evaluations by rating "Win/Tie/Lose" between the two samples. As shown in Table 8, the user simulators exhibit a notably superior performance compared to real users, particularly when it comes to the naturalness of responses in multi-turn conversations, which showcases the impressive language generation capabilities inherent in LLMs. Furthermore, even compared with human-annotated dialogues, the GPT3.5-based simulator shows competitive performance. These results validate the reliability of adopting GPT3.5 as the user simulator.
@@ A.4 Interactive Evaluation Protocol
During the evaluation, each dialogue agent must engage with these simulators Deng et al. (2023e). During interactions, the dialogue agent and user simulator alternate in employing strategies in their responses with the ultimate aim of maximizing their own self-interest. The interactions continues until the conversational goal is achieved or the maximum number of turns T (i.e., T is set to 10 for both tasks) is reached. To determine goal achievement, we utilize AI feedback to assess whether the task goal has been reached. Specifically, in price negotiation task, we employ a separate GPT3.5 (i.e., $LLM_{rwd}$) to assess whether both parties have reached a deal. We prompt $LLM_{rwd}$ to generate feedback for the binary question “Have they reached a deal?”. If the output of $LLM_{rwd}$ indicates that both parties have reached an agreement, we consider this as goal achievement. In charity persuasion task, we additionally prompt the user simulator to express his willingness to make a donation at the end of each turn. In particular, we query the user simulator "Would you be interested in donating to Save the Children?". If the feedback is positive, we regard this as goal achievement. Conversely, if the goal is not achieved, the interaction continues.
Due to the subjectivity of the planning outcome as well as the variance of the LLM-generated output, we follow a common practice Wang et al. (2022); Deng et al. (2023e) to alleviate these issues by sampling the decoded sequences l (i.e., l is set to 10 for both tasks) times.
@@ B.2 Baselines Implementation Details
ProCoT: we follow Deng et al. (2023d) and prompt LLM to analyze the dialogue status and plan next strategy, and then generate a response based on the planned strategy. We provide its prompt design in Table 17.
@@
PPDPP: we follow Deng et al. (2023e) and adopt the BERT [^7] https://huggingface.co/google-bert/bert-base-uncased model Devlin et al. (2018) as our external planner. We implement PPDPP based on the training details provided in the original paper. We have made adjustments to the task instructions and role-playing descriptions, adapting them for use in the context of charity persuasion.